OpenStudy (anonymous):

WILL FAN AND MEDAL!!!! Variable x is 7 more than variable y. Variable x is also 1 less than y. Which of the following pairs of equations best models the relationship between x and y? x = 1y x = y + 7 x = y + 7 x = y − 1 y = x + 8 y = x − 1 y = 7x y = x + 1

OpenStudy (texaschic101):

" is " means equals variable x is 7 more then variable y x = y + 7 variable x is 1 less then variable y x = y - 1

OpenStudy (anonymous):

so is it A

OpenStudy (texaschic101):

no...check again

OpenStudy (anonymous):

oh its C

OpenStudy (texaschic101):

lol....no answer is B......answers are : x = y + 7 and x = y - 1

OpenStudy (anonymous):

oh i thought it was C because it had 8 and thats 7 more than 1

OpenStudy (texaschic101):

your basically writing the equation like you " read " them x is (7 more then y) x = y + 7 x is (1 less then y) x = y - 1

OpenStudy (texaschic101):

it just takes practice....you will find that their is a " math language "

OpenStudy (anonymous):

oh ok thank you now i get it

OpenStudy (texaschic101):

glad to help :)